OpenStudy (anonymous):

If someone can help me through these trig identities that would be amazing...

OpenStudy (anonymous):

These are the questions i need to do..

OpenStudy (anonymous):

For a) they converted tan to one of its other ways of showing up, which is sin(x)/cos(x) and when you add those together, you get 1+sin(x)/cos(x) the denominator does not change because it is simple addition. and just about the same thing for the rest.. just look up identities.

OpenStudy (anonymous):

(sin x + cosx)^2 = (foil) sin^2x + 2sinxcosx + cos^2x = (sin^2x + cos^2x = 1) 1 = 2sinxcosx